If f(x)=3x+11, what is the value of (f^-1 o f)(2)....I don't even know where to start on this one!!!
Answer by venugopalramana(3286)

If f(x)=3x+11, what is the value of (f^-1 o f)(2)....I don't even know where to start on this one!!!
LET ME TELL YOU FEW THINGS FIRST...IF
F(X)=3X+11...THEN F(2) MEANS PUT 2 INPLACE OF X ...YOU GET F(2)=3*2+11=17
F^-1(X) MEANS LIKE THIS ..CALL Y=F(X)=3X+11 AS GIVEN ..THEN YOU HAVE TO FIND X INTERMS OF Y.AS WE HAVE 3X=Y-11
X=(Y-11)/3...THIS FUNCTION (Y-11)/3 IS CALLED F^-1(X) = F(Y)=(Y-11)/3
SO F(X) AND F^-1(X ARE INVERSE PROCESSES.SO IF YOU APPLY BOTH TO SOME X YOU GET BACK X..THAT IS F^-1 OF F(X)=X..LET US CJHECK WITH OUR EXAMPLE..
F(X)=3X+11..SO F^-1{3X+11)= {PUT 3X+11 IN PLACE OF Y SINCE F^-1(X)=F(Y)=(Y-11)/3
..SO WE GET (3X+11-11)/3=3X/3=X......

IN THE PRESENT PROBLEM.....
F(X)=3X+11....SO F(2)=3*2+11=17
F^-1(X)=F(Y)=(Y-11)/3....SO F^-1(F(2))=F^-1(17)=(17-11)/3=6/3=2...SO WE GET BACK OUR X=2 AS THE ANSWER AS PROVED ABOVE....